# Distribution Backtracking Builds A Faster Convergence Trajectory for One-step Diffusion Distillation

### v1

@@ B.1 Dataset Setup
The CelebA (Celebrities Attributes) dataset (Liu et al., 2015) is a large-scale collection of facial images used for facial attribute recognition. With over 200,000 images of celebrities, each image is annotated with 40 attributes, including features like hairstyle, gender, and age. In addition to attribute annotations, each image also includes annotations for facial landmarks such as eyes, nose, and mouth. For our experiments, we utilize the 64$\times$64 version of the CelebA dataset.
@@
The LSUN (Large Scale Scene Understanding) dataset (Yu et al., 2015) is a large-scale dataset for scene understanding in visual tasks within deep learning. Encompassing numerous indoor scene images, it spans various scenes and perspectives. The original LSUN dataset comprises multiple sub-datasets; in this study, we experiment on the LSUN Church and Bedroom sub-datasets with a resolution of 256$\times$256.
@@ B.2 Experiment Setup
When using the pre-trained EDM as the teacher model and the student generator, we use Adam optimizers to train the student generator $G$ and $s_{\phi}$, with both learning rates set to $1e^{-5}$. The training consisted of 500,000 epochs on four NVIDIA 3090 GPUs, and the batch size per GPU is set to 8. The training ratio between $s_{\phi}$ and $G$ remains at $1:1$. The pre-trained EDM models are all obtained from the official release of EDM (Karras et al., 2022) in https://github.com/NVlabs/edm. In the Degradation stage, we trained for 200 epochs total, saving a checkpoint every 50 epochs, resulting in a total of 5 intermediate nodes along the degradation path $\{s_{\theta_{i}}^{\prime}|i=0,1,2,3,4\}$. In the Distribution Backtracking stage, when $i\geq 3$, each checkpoint was trained for 1,000 steps. When $i<3$, each checkpoint was trained for 10,000 steps. The remaining steps were used to distill the original teacher model $s_{\theta_{0}}^{\prime}$.
Figure 7: Text-to-images generation samples by original LCM-LoRA and DisBack model distilled from LCM-LoRA with 512$\times$512 resolution.
@@
When using the pre-trained CD as the teacher model and the student generator, we set $\sigma_{max}$ to 80 and keep it constant during the single-step generation process. We use SGD and AdamW optimizers during training to train the generator $G$ and $s_{\phi}$, with learning rates set to $1e^{-3}$ and $1e^{-4}$, respectively. The training consisted of 10,000 epochs on one NVIDIA A100 GPU, and the batch size per GPU is set to 2. The training ratio between $s_{\phi}$ and $G$ remains at $4:1$. Since the original Consistency Model does not provide the pre-trained model on FFHQ64 and AFHQv264 datasets, we adopt the hyperparameters provided for the ImageNet64 dataset, setting the class conditional parameter to False and the number of training steps to 1,000,000 to train the Consistency Distillation model on the FFHQ64 and AFHQ64 dataset. In the Degradation stage, we trained for 200 epochs total and saved the checkpoint every 50 epochs, resulting in a total of 5 intermediate nodes along the degradation path $\{s_{\theta_{i}}^{\prime}|i=0,1,2,3,4\}$. In the Distribution Backtracking stage, when $i\geq 3$, each checkpoint was trained for 500 steps. When $i<3$, each checkpoint was trained for 1000 steps. The remaining steps were used to distill the original teacher model $s_{\theta_{0}}^{\prime}$.
@@
When distilling the SDXL model, the teacher model and the student generator are both initialed by the pre-trained SDXL model on the huggingface (model id is ‘stabilityai/stable-diffusion-xl-base-1.0’). We use Adam optimizers to train $G$ and $s_{\phi}$, with learning rates set to $1e^{-3}$ and $1e^{-2}$, respectively. The training consisted of 50,000 epochs on one NVIDIA A100 GPU, and the batch size per GPU is set to 1. The training ratio between $s_{\phi}$ and $G$ remains at $1:1$. The training prompts are obtained from LAION-Aesthetics. In the Degradation stage, we trained for 1,000 epochs total and saved the checkpoint every 100 epochs, resulting in a total of 10 intermediate nodes along the degradation path $\{s_{\theta_{i}}^{\prime}|i=0,1,...,9\}$. In the Distribution Backtracking stage, each checkpoint was trained for 1,000 steps. The remaining steps were used to distill the original teacher model $s_{\theta_{0}}^{\prime}$.
@@ C.1 KL Divergence of DisBack
$D_{\mathrm{KL}}(q\parallel p)=E_{q}\left[\log\frac{q}{p}\right]$ (10)
@@
$\nabla_{\eta}D_{\mathrm{KL}}\left(q^{G}_{t}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)=E_{t,\epsilon}\left[s_{\phi}\left(\bm{x}_{t},t\right)-s_{\theta}\left(\bm{x}_{t},t\right)\right]\frac{\delta\bm{x}_{t}}{\delta\eta}$ (12)
@@ C.2 Stable Target Field
Depending on the network prediction, the diffusion model can be divided into different types, including $\epsilon$ prediction (Karras et al., 2022) and $x_{0}$ prediction (Song et al., 2021a; Ho et al., 2020; Nichol & Dhariwal, 2021). When the score $\nabla_{\bm{x}_{t}}\log p_{t}(x_{t})$ is estimated by Eq.(18), it can be converted to $\epsilon$, $x_{0}$ and $v$ by a series of transformations.
$\hat{\epsilon}\approx-\sigma_{t}\nabla_{\bm{x}_{t}}\log p_{t}(x_{t})$ (19)
@@ Appendix D Discussion
D.1 Training Efficiency of DisBack
While DisBack involves an iterative optimization process during training, the optimization objective of $s_{\phi}\left(\bm{x}_{t},t\right)$ aims to minimize the loss of the standard diffusion model based on Eq.(5), and the objective of student generator aims to minimize the KL divergence in Eq.(6). These two optimization processes do not entail an adversarial training as in GANs. Consequently, the optimization process tends to be more stable. A recent work Monoflow (Yi et al., 2023) also discusses in GANs training a vector field is obtained to guide the optimization of the generator, but the vector field derives from the discriminator and the instability is not mitigated.
$\min_{\phi}E_{t,\epsilon}\left\|s_{\phi}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ (21)
@@ D.1 Training Efficiency of DisBack
D.2 Vector Field
In our research, each of the estimated score functions $s_{\theta_{i}}$, for $i$ ranging from 0 to $N$, delineates a vector field $\mathbb{R}^{3\times W\times H}\mapsto\mathbb{R}^{3\times W\times H}$. We make a strong assumption behind our proposed method that these score functions represent existing or non-existent distributions and that they altogether imply a transformation path between $s_{\theta}$ and the student generator $G_{0}$. Nevertheless, a score fundamentally constitutes a gradient field, signifying the gradient of the inherent probability density. A vector field is a gradient field when several conditions are satisfied, including path independence, continuous partial derivatives, and zero curls (Matthews, 1998). The vector field, as characterized by the score functions, may not meet these conditions, and thus there is not a potential function or a probability density function. Such deficiencies could potentially hinder the successful training of the student generator and introduce unforeseen difficulties in the distillation process. Specifically, in instances where $s_{\theta}$ does not precisely represent a gradient field, a highly probable scenario considering $s_{\theta}$ is a neural network, the samples generated from $s_{\theta}$ could encompass failure cases. Although our empirical studies exemplify the effectiveness of the proposed Generative Backtracking Model, the detrimental effects of the discussed issue remain unclear. We will further explore this issue in our future work.
@@ Appendix E Additional Details in Pre-experiments
E.1 A Toy Experiment on Gaussian Mixture Distribution
To validate the feasibility of the proposed DisBack, we conduct experiments on two-dimensional Gaussian mixture data. First, we randomly select $10$ Gaussian distributions mixed as the training distribution $q_{0}$. Next, we construct a ResNet MLP as the two-dimensional diffusion model $s_{\theta}$ and train it using the created mixture Gaussian distribution. Similarly, we construct a simple MLP as the student generator $G_{stu}$ and train a model $s_{\phi}$ with the same architecture as $s_{\theta}$ using generated data. Therefore, we can use $s_{\theta}$ and $s_{\phi}$ to train the student generator $G_{stu}$. During the training process, we visualize the distribution of the student generator and training data to intuitively demonstrate the changes in the student generator distribution under the proposed training framework. The distribution of $G_{stu}$ during the training process is shown in Figure 8. As training progresses, the generated distribution $q^{G}$ initially expands outward and then gradually convergents towards the training distribution. The results show that the proposed method for training the student generator in the paper is effective.
@@ E.1 A Toy Experiment on Gaussian Mixture Distribution
E.2 Gradient Orientation Verification of DisBack
As mentioned in Section 3.1 of the main paper, when updating $G_{stu}$ using Eq.(6), $s_{\theta}\left(\bm{x}_{t},t\right)$ provides a gradient towards the training distribution, while $s_{\phi}\left(\bm{x}_{t},t\right)$ provides a gradient toward the generated distribution.
@@ F.1 Samples of Image Generation
F.2 Samples Text-to-Image generation
Appendix G Ethical Statement
G.1 Ethical Impact
@@
G.2 Notification to Human Subjects
@@
Dear volunteers, we would like to thank you for supporting our study. We propose the Generative Backtracking Model, which uses a backtracking strategy to achieve stable and fast distillation and high-quality single-step generation.
@@
Figure 10: Additional Samples form conditional FFHQ64.
Figure 11: Additional Samples form conditional AFHQv264.
Figure 12: Additional Samples form conditional ImageNet64.

### v3

@@ B.1 Dataset Setup
The LSUN (Large Scale Scene Understanding) dataset (Yu et al., 2015) is a large-scale dataset for scene understanding in visual tasks within deep learning. Encompassing numerous indoor scene images, it spans various scenes and perspectives. The LSUN dataset comprises multiple sub-datasets, in this study, we use the LSUN Cat and Bedroom sub-datasets with a resolution of 256$\times$256.
Figure 8: Samples from the initial generator, intermediate teacher trajectory nodes. Here $s_{\theta_{0}}^{\prime}$ is the teacher model. The teacher model is the pre-trained EDM model on the AFHQv2 dataset, the student generator is FastGAN.
Figure 9: Samples from the initial generator, intermediate teacher trajectory nodes. Here $s_{\theta_{0}}^{\prime}$ is the teacher model. The teacher model and the student generator are both the pre-trained EDM model on the ImageNet dataset.
@@ B.2 Experiment Setup
For experiments on FFHQ 64x64, AFHQv2 64x64, and ImageNet 64x64 datasets, the pre-trained models are provided by the official release of EDM Karras et al. (2022). We use Adam optimizers to train the student generator $G$ and $s_{\phi}$, with both learning rates set to $1e^{-5}$. The training consisted of 50,000 iterations on four NVIDIA 3090 GPUs, and the batch size per GPU is set to 8. The training ratio between $s_{\phi}$ and $G$ remains at $1:1$. In the Degradation stage, we trained for 200 epochs total, saving a checkpoint every 50 epochs, resulting in a total of 5 intermediate nodes along the degradation path $\{s_{\theta_{i}}^{\prime}|i=0,1,2,3,4\}$. In the Distribution Backtracking stage, when $i\geq 3$, each checkpoint was trained for 1,000 steps. When $i<3$, each checkpoint was trained for 10,000 steps. The remaining steps were used to distill the original teacher model $s_{\theta_{0}}^{\prime}$.
@@
For experiments on LSUN bedroom and LSUN cat datasets, the pre-trained EDM models are provided by the official release of Consistency Model Song et al. (2023). During the training, we set $\sigma_{max}$ to 80 and keep it constant during the single-step generation process. We use SGD and AdamW optimizers during training to train the generator $G$ and $s_{\phi}$, with learning rates set to $1e^{-3}$ and $1e^{-4}$, respectively. The training consisted of 10,000 iterations on one NVIDIA A100 GPU, and the batch size per GPU is set to 2. The training ratio between $s_{\phi}$ and $G$ remains at $4:1$. In the Degradation stage, we trained for 200 epochs total and saved the checkpoint every 50 epochs, resulting in a total of 5 intermediate nodes along the degradation path $\{s_{\theta_{i}}^{\prime}|i=0,1,2,3,4\}$. In the Distribution Backtracking stage, when $i\geq 3$, each checkpoint was trained for 500 steps. When $i<3$, each checkpoint was trained for 1000 steps. The remaining steps were used to distill the original teacher model $s_{\theta_{0}}^{\prime}$.
@@
When distilling the SDXL model, the teacher model and the student generator are both initialed by the pre-trained SDXL model on the huggingface (model id is ‘stabilityai/stable-diffusion-xl-base-1.0’). We use Adam optimizers to train $G$ and $s_{\phi}$, with learning rates set to $1e^{-3}$ and $1e^{-2}$, respectively. The training consisted of 50,000 iterations on one NVIDIA A100 GPU, and the batch size per GPU is set to 1. The training ratio between $s_{\phi}$ and $G$ remains at $1:1$. The training prompts are obtained from LAION-Aesthetics. In the Degradation stage, we trained for 1,000 epochs total and saved the checkpoint every 100 epochs, resulting in a total of 10 intermediate nodes along the degradation path $\{s_{\theta_{i}}^{\prime}|i=0,1,...,9\}$. In the Distribution Backtracking stage, each checkpoint was trained for 1,000 steps. The remaining steps were used to distill the original teacher model $s_{\theta_{0}}^{\prime}$.
@@ C.1 KL Divergence of DisBack
$D_{\mathrm{KL}}(q\parallel p)=\mathbb{E}_{q}\left[\log\frac{q}{p}\right]$ (10)
@@
$\nabla_{\eta}D_{\mathrm{KL}}\left(q^{G}_{t}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)=\mathbb{E}_{t,\bm{\epsilon}}\left[s_{\phi}\left(\bm{x}_{t},t\right)-s_{\theta}\left(\bm{x}_{t},t\right)\right]\frac{\delta\bm{x}_{t}}{\delta\eta}$ (12)
@@ C.2 Stable Target Field
Depending on the network prediction, the diffusion model can be divided into different types, including $\bm{\epsilon}$ prediction (Karras et al., 2022) and $x_{0}$ prediction (Song et al., 2021a; Ho et al., 2020; Nichol & Dhariwal, 2021). When the score $\nabla_{\bm{x}_{t}}\log p_{t}(x_{t})$ is estimated by Eq.(18), it can be converted to $\bm{\epsilon}$, $x_{0}$ and $v$ by a series of transformations.
$\hat{\bm{\epsilon}}\approx-\sigma_{t}\nabla_{\bm{x}_{t}}\log p_{t}(x_{t})$ (19)
@@ Appendix D Discussion
D.1 The gradient orientation of score distillation
Figure 10: An illustration of the generator update by the minimization in Eq. (6).
To better illustrate the process of training a generator with score distillation, we provide a more intuitive explanation. As shown in Figure 10, given a noisy data $\bm{x}_{t}$, the direction of $s_{\phi}\left(\bm{x}_{t},t\right)$ points to the generated distribution and the direction of $s_{\theta}\left(\bm{x}_{t},t\right)$ points to the training distribution. In Sec.E.3, we used 2D data to illustrate the gradient directions of $\bm{x}_{t}$ on the teacher model $s_{\theta}$ and the auxiliary diffusion model $s_{\phi}$. Thus, the direction of $s_{\theta}\left(\bm{x}_{t},t\right)-s_{\phi}\left(\bm{x}_{t},t\right)$ is from the generated distribution towards the training distribution. Ideally, if the predicted score of $s_{\phi}$ and $s_{\theta}$ are both correct, this gradient leads $G$ to update in the right direction.
D.2 Two types of mismatch issue
We find that there are two types of mismatch issues during the training process: (1) The mismatch issue between the initially generated samples and $s_{\phi}$. (2) The mismatch issue between the initially generated samples and $s_{\theta}$. For the first type of mismatch issue, it prevents $s_{\phi}$ from accurately describing the distribution of the generator, leading to unreliable network predictions and introducing biases when computing the generator’s gradients. The combined effect of these two mismatch issues exacerbates the negative impact on the optimization of the generator. Fortunately, among the three initialization methods of $s_{\phi}$ proposed in Section 3, when $s_{\phi}$ is initialized based on the generated images, its distribution becomes close to the initially generated distribution, which resolves the first type of mismatch issue. In this paper, we directly use the degraded teacher model $s_{\theta_{N}}^{\prime}$ to initialize $s_{\phi}$. Because the degraded teacher model’s distribution is closer to the initial generator’s distribution, $s_{\phi}$ initialized by $s_{\theta_{N}}^{\prime}$ can approximate the initial generation distribution and accurately predict the scores of the generated samples from the beginning. However, since $s_{\theta}$ remains fixed during the training, the second type of mismatch issue cannot be avoided through initialization. Consequently, in this paper, we mainly focus on discussing the second type of mismatch issue.
D.3 Training Efficiency of DisBack
While DisBack involves an iterative optimization process during training, the optimization objective of $s_{\phi}\left(\bm{x}_{t},t\right)$ aims to minimize the loss of the standard diffusion model based on Eq.(21), and the objective of student generator aims to minimize the KL divergence in Eq.(22). These two optimization processes do not entail adversarial training as in GANs. Consequently, the optimization process tends to be more stable. A recent work Monoflow (Yi et al., 2023) also discusses in GANs training a vector field is obtained to guide the optimization of the generator, but the vector field derives from the discriminator and the instability is not mitigated.
$\min_{\phi}\mathbb{E}_{t,\bm{\epsilon}}\left\|s_{\phi}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ (21)
$\min_{\eta}\mathbb{E}_{t,\bm{\epsilon}}D_{KL}\left(q_{t}^{G}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)$ (22)
@@ D.3 Training Efficiency of DisBack
D.4 Vector Field
In our research, each of the estimated score functions $s_{\theta_{i}}^{\prime}$, for $i$ ranging from 0 to $N$, delineates a vector field $\mathbb{R}^{3\times W\times H}\mapsto\mathbb{R}^{3\times W\times H}$. We make a strong assumption behind our proposed method that these score functions represent existing or non-existent distributions and that they altogether imply a transformation path between $s_{\theta}$ and the student generator $G^{0}_{stu}$. Nevertheless, a score fundamentally constitutes a gradient field, signifying the gradient of the inherent probability density. A vector field is a gradient field when several conditions are satisfied, including path independence, continuous partial derivatives, and zero curls (Matthews, 1998). The vector field, as characterized by the score functions, may not meet these conditions, and thus there is not a potential function or a probability density function. Such deficiencies could potentially hinder the successful training of the student generator and introduce unforeseen difficulties in the distillation process. Specifically, in instances where $s_{\theta}$ does not precisely represent a gradient field, a highly probable scenario considering $s_{\theta}$ is a neural network, the samples generated from $s_{\theta}$ could encompass failure cases. Although our empirical studies exemplify the effectiveness of the proposed DisBack, the detrimental effects of the discussed issue remain unclear. We will further explore this issue in our future work.
@@ Appendix E Additional Details in Pre-experiments
E.1 Distribution mismatch issues
Figure 11: The results of the pre-experiments on mismatch degree. (a) $s_{\theta}$ and $G$ are both initialized by the pre-trained EDM (Karras et al., 2022) on FFHQ. (b) $s_{\theta}$ and $G$ are both initialized by the pre-trained EDM on ImageNet. The mismatch degree on the generated data is greater than on the training data, especially when the noise scale is low.
Before conducting our research, we first carried out preliminary experiments to demonstrate that the proposed mismatch issue does indeed exist when using the endpoints of pre-trained diffusion models as teacher models. Using the method proposed in Eq. 9, we conducted experiments with the pre-trained EDM model on the ImageNet and FFHQ datasets. We calculated the mismatch degree separately on the student model’s initial generated data and the teacher model’s original training data, and the results are shown in Fig. 11. We can see that, on both datasets, the mismatch degree on the generated data of the pre-trained model is greater than that on the real training data, especially when the noise scale is small. This aligns with our hypothesis stated in Sec. 1, demonstrating that directly using the endpoint of a pre-trained model as the teacher model leads to a distribution mismatch problem and causes the unreliable predictions of the teacher model.
@@ E.1 Distribution mismatch issues
E.2 A Toy Experiment on Gaussian Mixture Distribution
To validate the feasibility of the proposed DisBack, we conduct experiments on two-dimensional Gaussian mixture data. First, we randomly select $10$ Gaussian distributions mixed as the training distribution $q_{0}$. Next, we construct a ResNet MLP as the two-dimensional diffusion model $s_{\theta}$ and train it using the created mixture Gaussian distribution. Similarly, we construct a simple MLP as the student generator $G_{stu}$ and train a model $s_{\phi}$ with the same architecture as $s_{\theta}$ using generated data. Therefore, we can use $s_{\theta}$ and $s_{\phi}$ to train the student generator $G_{stu}$. During the training process, we visualize the distribution of the student generator and training data to intuitively demonstrate the changes in the student generator distribution under the proposed training framework. The distribution of $G_{stu}$ during the training process is shown in Figure 12. As training progresses, the generated distribution $q^{G}$ initially expands outward and then gradually convergents towards the training distribution. The results show that the proposed method for training the student generator is effective.
E.3 Gradient Orientation Verification of DisBack
As mentioned in Sec. 3, when updating $G_{stu}$ using Eq.(6), $s_{\theta}\left(\bm{x}_{t},t\right)$ provides a gradient towards the training distribution, while $s_{\phi}\left(\bm{x}_{t},t\right)$ provides a gradient toward the generated distribution.
@@ Appendix F Additional Samples from DisBack
Appendix G Additional Ablation Study
We further explored the impact of different numbers of checkpoints along the degradation path and various degradation strategies on the performance of DisBack on the ImageNet64 dataset. The results are shown in Tab.6. When degrading for $200iterations$ under default settings, the FID score with $N=3$ intermediate checkpoints was $5.23$, which outperformed the original Diff-instruct’s FID of $5.57$. The optimal FID of $1.38$ was achieved with $N=5$.
Table 6: The conditional generation performance of DisBack on ImageNet 64x64 dataset.
DisBack FID ($\downarrow$)
N=1 5.96
N=3 4.88
N=5 1.38
N=11 9.15
N=11(Degradation iteration =400) 244.72
However, with $N=11$, the FID score dropped to $8.35$. When there are too many checkpoints in the path, the distributions of the checkpoints (e.g., $s_{\theta_{1}0}^{\prime},s_{\theta_{1}9}^{\prime},s_{\theta_{8}}^{\prime}$) are very close to the distribution of the initial generator. Training with these checkpoints provides limited progress for the generator. Furthermore, having too many checkpoints complicates the checkpoint transition scheduler during distillation, making it difficult to manage effectively. This often leads to inefficient updates for the generator across many iterations, wasting time without meaningful improvement. When the degradation iteration count is set to 400, DisBack fails to correctly distill the student generator. This is because excessive degradation iterations result in checkpoints near the initial generator on the degradation path being unable to generate sensible samples. Using these checkpoints for distillation provides no reasonable or effective guidance to the generator, ultimately causing it to fail to learn any meaningful information. In summary, when obtaining the degradation path, it is essential not to overly degrade the teacher model. The degraded teacher model’s distribution should be close to the initial generator’s distribution while still being able to produce relatively reasonable samples. Additionally, having too many or too few checkpoints along the path can adversely affect the final performance.
Appendix H Failure Examples
Fig. 14 presents several failure cases of DisBack.
In terms of FFHQ, AFHQv2, and ImageNet, while these images already capture the features of the corresponding datasets, the generated results lack accurate and clear backgrounds. The potential reasons for this include the fact that these datasets primarily focus on learning foreground content, with low requirements for image backgrounds, making the model difficult to clear backgrounds.
As for LSUN Cat and Bedroom, DisBack successfully generates details such as the cat’s fur and the bed’s texture, but it does not generate the overall shape and the detailed structure. This may be because the model does not capture the overall information of the data, only capturing local content. This issue may stem from the inherent limitations of U-Net, resulting in poor generation of overall structures in rare cases.
In the future, attempts will be made to use more advanced teacher models or improve the distillation algorithm to overcome these limitations. Moreover, we will further explore more advanced generator architectures such as StyleGAN Karras et al. (2020; 2021) to achieve higher-quality generation.
Figure 14: Failure examples.
Appendix I Ethical Statement
I.1 Ethical Impact
@@
I.2 Notification to Human Subjects
@@
Dear volunteers, we would like to thank you for supporting our study. We propose the Distribution Backtracking Distillation, which introduces the convergence trajectory into the score distillation process to achieve efficient and fast distillation and high-quality single-step generation.
@@
Figure 15: Additional Samples form conditional FFHQ 64x64.
Figure 16: Additional Samples form conditional AFHQv2 64x64.
Figure 17: Additional Samples form conditional ImageNet 64x64.